# Deep Learning of Multivariate Extremes via a Geometric Representation

### v1

@@ 5.2.1 Joint distribution of hs, ws, and mslp: $d=3$
Figure 8: The estimated unit-level set (left) and the extended ADF (right) for hs, ws, and mslp at location 01 of the metocean data. All data are plotted on standard Laplace margins.
Here we present results for the three metocean variables observed at location 01; results for the remaining two locations are presented in Appendix C. Figure 8 plots the estimates of the three-dimensional unit-level set and the extended ADF for the three variables, with the sample clouds transformed to standard Laplace margins. We can see that the estimated unit-level sets capture well the shape and joint tail behaviour of the sample cloud. We also plot bivariate unit-level set slices for each of the three pairs of variables in Figure 9. For these, we set $\epsilon=0.01$ (see Section 3.5), resulting in a reasonable amount of validation data for each unit-level set slice. We observe that the estimated slices closely match the shapes of the corresponding bivariate sample clouds, again suggesting good model fits to these data. Corresponding plots for the remaining two locations are provided in Figures A6A9.
Goodness-of-fit is further verified using the truncated gamma QQ plot, extended ADF diagnostic, and return level set probabilities, as described in Section 3.5. The plots, provided in Figure A3 of Appendix C, illustrate strong agreement between model and empirical quantiles across all three diagnostics. The goodness-of-fit analyses for the remaining two locations also indicate good model fits over all diagnostics (Figures A4A5 of Appendix C). For all three locations, we also provide animated three dimensional plots for the unit-level sets and extended ADFs in the Supplementary Material.
Figure 9: Scaled bivariate sample clouds with $\epsilon=0.01$ for pairwise combinations of ws, hs, and mslp at location 01, transformed onto common Laplace margins; the red shapes outline the corresponding estimated bivariate unit-level set slices.
@@ 5.2.2 Joint distribution of hs along transect: $d=5$
For the $d=5$ joint distribution of significant wave height hs, we cannot visualise the unit-level set estimate. However, there are 10 possible bivariate unit-level slices corresponding to different pairs of locations; three of these are illustrated in Figure 10. In this case, we set $\epsilon=0.05$ to ensure a reasonable amount of validation data in each slice, accounting for the fact angular observations are more sparse in higher dimensions (relative to $d=3$). As previously, we observed that the estimated slices closely match the shapes of the corresponding bivariate sample cloud, indicating that the DeepGauge model can capture extremal dependence of hs across all five locations. Goodness-of-fit diagnostics, as illustrated in Figure A10 of Appendix C, show generally good agreement between model and empirical quantiles, albeit with some small divergences at extremely high quantiles (for the truncated gamma QQ plot).
Figure 10: Scaled bivariate sample clouds with $\epsilon=0.05$ for pairwise combinations of hs at three locations, transformed into common Laplace margins; the red shapes outline the corresponding estimated bivariate unit-level set slices.
The DeepGauge framework provides fantastic model fits for our metocean data, in both $d=3$ and $d=5$ dimensions. The estimated unit-level sets facilitate inference of the entire extremal dependence structure of the data, not only dependence in a single orthant of $\mathbb{R}^{d}$, without the need for restrictive parametric models or low-order summary measures (see Section 1). Whilst we have omitted this part of the study for brevity, we note that estimates of our novel extended ADF permit estimation of joint tail probabilities on arbitrary hyper-cubes; our accompanying diagnostics for the ADF estimates suggest high accuracy of such estimates. Finally, we remark that our favourable results for the case where $d=5$ are particularly encouraging. Flexible models for extremal dependence are typically limited to $d\leq 3$, but we have shown that this limitation does not hold for the DeepGauge framework.
@@ 6 Discussion
We introduce novel theoretical contributions for the geometric representation of multivariate extremes, which include extensions of the Wadsworth and Tawn, (2013) model and a demonstration that many of the results introduced by Nolde and Wadsworth, (2022) can be extended to data on Laplace margins. We further introduce the DeepGauge modelling framework and, through rigorous theoretical treatment and in contrast to all existing models, we prove that our framework gives estimates that satisfy all of the required theoretical properties of unit-level sets. In practice, this results in consistent inference of extremal dependence and provides a means to estimate joint tail properties; we demonstrate this by evaluating the performance of the DeepGauge framework on simulated and observed data in Sections 4 and 5, respectively. Unlike the majority of existing models for multivariate extremes, the DeepGauge framework is not limited to asymptotically dependent data nor low dimension $d\leq 3$. Thus, our approach represents a significant step towards flexible, non-parametric, and scalable models for multivariate extremes.
@@
One noticeable omission from the DeepGauge framework is a model for the angular density $f_{\bm{W}}(\bm{w}),\bm{w}\in\mathcal{S}^{d-1}$; see Section 3.4 for further discussion. Several existing approaches propose non-parametric techniques for this estimation, albeit in low dimensional ($d\leq 3$) settings (Papastathopoulos et al.,, 2024; Murphy-Barltrop et al., 2024a, ; Simpson and Tawn, 2024b, ). Models for the angular density offer the advantage of simulation from the model described in equation (7), from which one can obtain probability estimates in joint tail regions of any form, i.e., more general regions than those illustrated in Figure 4; see Wadsworth and Campbell, (2024) for instance. The estimation of, and simulation from, the angular density function $f_{\bm{W}}(\bm{w})$ therefore represents an important area for future work in the context of multivariate extremes.
@@
Acknowledgements
Reetam Majumder was supported by grants from the United States Geological Survey’s National Climate Adaptation Science Center (G24AC00197), and the National Science Foundation (DMS2152887). The authors would like to thank Phil Jonathan for providing the data, Ryan Campbell for access to code, and Lambert de Monte for helpful discussions.
@@ A.1 Neural network representation
We here describe the construction of $r_{\tau}(\cdot)$ and $g(\cdot)$ used multi-layer perceptrons. Let $r_{\tau}(\bm{w}):=\exp\{m_{\bm{\psi}}(\bm{w})\},$ where $m_{\bm{\psi}}(\bm{w})$ is a neural network, parameterised by $\bm{\psi}$, that comprises $N\in\mathbb{N}$ hidden layers. Then $m_{\bm{\psi}}$ is a composition of the form ${m_{\bm{\psi}}(\cdot)=m_{\bm{\psi}}^{(N+1)}\circ\dots\circ\mathbf{m}^{(1)}_{\bm{\psi}}(\cdot)}$, where, for $j=1,\dots,N$, the output from the $j$-th hidden layer, $\mathbf{m}^{(j)}_{\bm{\psi}}:\mathbb{R}^{n_{j-1}}\mapsto\mathbb{R}^{n_{j}}$, can be written as
@@ A.2 Training and pre-training
The minimisation in (8) is performed using the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba,, 2014), which is a variant of mini-batch stochastic gradient descent [^1] We use the default settings of the hyper-parameters for Adam.. We utilise the R interface to keras (Allaire and Chollet,, 2021). Prior to training, we uniformly-at-random partition the data into $80\%$ training and $20\%$ validation, with this partition consistent across the training stage for both $r_{\tau}(\bm{w})$ and $\tilde{g}(\bm{w})$. To mitigate overfitting of the neural network model, we train with checkpoints and early-stopping (Prechelt,, 2002). Neural network models are optimised for a finite number of pre-specified iterations, say $M$, by minimising the loss function evaluated on the training data. The validation data are not used to optimise the neural network parameters. Instead, at every iteration of the training algorithm, we evaluate the loss function on the validation data. Then, the “best fitting” model is determined to be that which which minimises the loss function, evaluated on the validation data, across all epochs. Early-stopping ensures that the training scheme does not necessarily run for all $M$ iterations; training stops early if the validation loss has not decreased in the last $\Delta\in\mathbb{N}$ iterations. We also add $L_{1}$ and $L_{2}$ penalties to the estimable neural network parameters, $\psi$, to provide further model regularisation; throughout we adopt these penalties with equal shrinkage weight of $1\times 10^{-4}$. For further details on fitting and regularising deep neural network models, see Goodfellow et al., (2016).
@@
Typically, without prior knowledge of the function that one seeks to approximate, the estimable parameter set, $\bm{\psi},$ of a neural network model is randomly initialised prior to training; thus when training our neural network model for the exceedance threshold $r_{\tau}(\cdot)$ we always choose to take this approach, as we have no prior knowledge of its structure. However, for training of the neural network that comprises the gauge function $g(\cdot)$, we can exploit some of the theoretical structure of gauge functions to perform pre-training, which is a popular technique used in applications of neural networks to reduce computation times. Pre-training use parameter estimates of a neural network trained for one task as initial parameter estimates for a neural network model designed for a different task (with equivalent architecture; see Goodfellow et al., (2016), Ch. 8). Following Corollary 2.2, we can use an estimate of the radial exceedance threshold, say $\hat{r}_{\tau}(\bm{w}),$ to obtain an initial estimate for the gauge function, say $\hat{g}_{\tau}(\bm{w})$; this is achieved by replacing the radial function $h(\cdot)$ in Corollary 2.2 with $\hat{r}_{\tau}(\cdot)$ and deriving the corresponding rescaled gauge function. In the two dimensional setting, Wadsworth and Campbell, (2024) use a similar idea to get an initial estimate for the unit-level set $\partial\mathcal{G}$. The authors also show that, for $\tau$ close to 1, the rescaled quantile set is a valid approximation of $\partial\mathcal{G}$. We use this idea to pre-train the neural network of the gauge function model $\tilde{g}(\cdot)$; we perform an initial optimisation of its weights subject to minimisation of the squared-error loss between the initial gauge function $\hat{g}_{\tau}(\bm{w})$ and the model output, $\tilde{g}(\bm{w})$. For $\alpha$, we always set its initial estimate to $d$, as this is the theoretical value attained by gauge functions for many popular copula models (Wadsworth and Campbell,, 2024). In unreported experiments, we found that this pre-training helps to mitigate numerical instability during training and improves accuracy of gauge function estimation.
@@ Appendix B Additional proofs
is standard Laplace distributed, where the $\log(2)$ here denotes the median of $X_{E}$. Given $\bm{w}\in\mathcal{S}^{d-1}_{+}$ we have, by definition, $g_{E}(\bm{w})=\lim_{t\to\infty}(-\log f_{\bm{X}_{E}}(t\bm{w})/t)$, where $f_{\bm{X}_{E}}(\cdot)$ denotes the continuous joint density function for $\bm{X}_{E}$. For any large $t$ satisfying $t>\max_{i=1,\dots,d}\{\log(2)/w_{i}\}>0$, we have that
@@
Given any large $u$ satisfying $u>\max_{i=1,\dots,d}\{\log(2)/w_{i}\}>0$, and applying similar reasoning as in the proof of Proposition 2.1 (Appendix B.1), we have that
@@
Given $\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$, set $\bm{c}:=\text{sgn}(\bm{w})=(\text{sgn}(w_{1}),\ldots,\text{sgn}(w_{d}))^{T}$, where $\text{sgn}(\cdot)$ denotes the regular signum function. Then, we have that
@@
By Proposition 2.5, $g_{\bm{c}\bm{X}}(\bm{w}^{*})=g(\bm{w}^{*}/\bm{c})$, with $\bm{w}^{*}/\bm{c}:=(w^{*}_{1}/c_{1},\ldots,w^{*}_{d}/c_{d})\in\mathcal{S}^{d-1}\setminus\mathcal{A}$ and $\text{sgn}(\bm{w})=\text{sgn}(\bm{w}^{*}/\bm{c})$. Thus, we have
@@
$\frac{\tilde{\mathfrak{r}}_{\bm{w}}}{||\bm{w}||_{\infty}}\geq\min_{i=1,\dots,d}\left\{\frac{w_{i}}{g(\bm{w})w_{i}}\right\}=\frac{1}{g(\bm{w})}.$
Recalling from Proposition 2.8 that $||\bm{w}||_{\infty}\Lambda(\bm{w})^{-1}=\tilde{\mathfrak{r}}_{\bm{w}}$, the first inequality follows. The second inequality follows directly from Proposition 2.8.

### v2

@@ 5.2.1 Joint distribution of hs, ws, and mslp: $d=3$
Figure 8: The estimated unit-level (left) and extended ADF (right) sets for hs, ws, and mslp at location 01 (on standard Laplace margins).
Here we present results for the three metocean variables observed at location 01; results for the remaining two locations are presented in Appendix C. Figure 8 plots the estimates of the three-dimensional unit-level set and the extended ADF for the three variables, with the sample clouds transformed to standard Laplace margins. We observe that the estimated unit-level set captures well the shape of the scaled sample cloud. We also plot bivariate unit-level set slices for each of the three pairs of variables in Figure 9. For these, we set $\epsilon=0.01$ (see Section 3.5), resulting in a reasonable amount of validation data for each unit-level set slice. We observe that the estimated slices closely match the shapes of the corresponding bivariate sample clouds, again indicating accurate model fits. Corresponding plots for the remaining two locations are provided in Figures A6A9 (Appendix C).
Goodness-of-fit is further verified using the truncated gamma QQ plot, the extended ADF diagnostic evaluated at $\bm{w}=\bm{1}_{d}/||\bm{1}_{d}||$, and the estimated return level set probabilities; see Section 3.5. The plots, provided in Figure A3 of Appendix C, illustrate strong agreement between model and empirical quantiles across all three diagnostics. The goodness-of-fit analyses for the remaining two locations also indicate good model fits over all diagnostics (Figures A4A5 of Appendix C). For all locations, we also provide animated three dimensional plots for the unit-level sets and extended ADFs in the Supplementary Material.
Figure 9: Scaled bivariate sample clouds with $\epsilon=0.01$ for pairwise combinations of ws, hs, and mslp at location 01 on standard Laplace margins; the red lines describe the estimated bivariate unit-level set slices.
Finally, to demonstrate a practical use case of the three dimensional model fits, we use the estimated extended ADFs, obtained from the estimated unit-level sets, to estimate probabilities for joint tail regions. These tail regions, or hypercubes, are illustrated in Figure 10 for metocean data of location $01$ (on Laplace margins), noting that the same hypercubes are considered across all three locations. The orange, blue and purple hypercubes correspond to the tail regions $\{\texttt{hs}_{L}>u_{0.99},\texttt{ws}_{L}>u_{0.99},\texttt{mslp}_{L}<-u_{0.99}\}$, $\{\texttt{hs}_{L}>u_{0.99},\texttt{ws}_{L}<-u_{0.99},\texttt{mslp}_{L}<-u_{0.99}\}$, and $\{\texttt{hs}_{L}<-u_{0.99},\texttt{ws}_{L}>u_{0.99},\texttt{mslp}_{L}<-u_{0.99}\}$ respectively, where $u_{0.99}$ denotes the 0.99 quantile of the standard Laplace distribution and $\cdot_{L}$ denote each metocean variable on Laplace margins. As noted previously, high wind speeds and wave heights combined with low sea level pressure are typical conditions for the most extreme storm events (Ewans and Jonathan, 2014b, ; Lobeto et al.,, 2024); this corresponds with the orange hypercube in Figure 10. Conversely, owing to the underlying physics of metocean phenomena, one would expect to observe very little data in the remaining hypercubes. Bootstrapped probability estimates for each hypercube and location are given in Appendix C; the results agree with what one would expect, in the sense that the estimates for the orange hypercubes are many orders of magnitude greater than the blue and purple regions. For example, at location $01$, the bootstrapped probability estimates for the orange region exist within the interval $[2.9\times 10^{-4},8.2\times 10^{-4}]$ (i.e., rare but feasible within a reasonable time window), whereas the corresponding estimates for the remaining regions are all smaller than $2.73\times 10^{-8}$. This indicates the estimated probabilities are accurately representing the observed tail structure.
Figure 10: Hypercube regions considered for the three metocean variables.
@@ 5.2.2 Joint distribution of hs along transect: $d=5$
For the $d=5$ joint distribution of significant wave height hs, we cannot visualise the unit-level set estimate. However, there are 10 possible bivariate unit-level slices corresponding to different pairs of locations; three of these are illustrated in Figure 11. We again set $\epsilon=0.01$ as this results in a reasonable amount of validation data for each slice. As previously, we observe that the estimated slices closely match the shapes of the corresponding bivariate sample cloud, indicating that the DeepGauge model can capture extremal dependence of hs across all five locations. Goodness-of-fit diagnostics, as illustrated in Figure A13 of Appendix C, show generally good agreement between model and empirical quantiles, albeit with slightly more divergence at extremely high quantiles relatively to the $d=3$ cases. This is especially true for the ADF diagnostic; however, this is not surprising considering the fact observations in each orthant are more sparse in higher dimensions (relative to $d=3$), and more data is generally required in such cases to ensure convergence to asymptotic forms.
We also test the DeepGauge framework on a transect with $d=8$ locations; these results are detailed in the Supplementary Material. For this case, diagnostics again indicate generally good performance, albeit with a slightly reduction in quality compared to the lower dimensional cases.
Figure 11: Scaled bivariate sample clouds with $\epsilon=0.01$ for three pairwise combinations of the hs variable over the five locations on standard Laplace margins; the red lines describe the estimated bivariate unit-level set slices.
Overall, the proposed diagnostics indicate that the DeepGauge framework provides accurate model fits for metocean data sets of $d=3$, $d=5$, and $d=8$ dimensions. The estimated unit-level sets facilitate inference of the entire extremal dependence structure of the data, not only dependence in a single orthant of $\mathbb{R}^{d}$, without the need for restrictive parametric models or low-order summary measures (see Section 1). We also remark that our results for the cases of $d=5$ and $d=8$ are particularly encouraging: flexible models for extremal dependence are typically limited to $d\leq 3$, but we have shown that this limitation does not hold for the DeepGauge framework.
@@ 6 Discussion
We have introduced novel theoretical contributions for the geometric representation of multivariate extremes, which include extensions of the Wadsworth and Tawn, (2013) model and a demonstration that many of the results introduced by Nolde and Wadsworth, (2022) can be extended to data on Laplace margins. We further proposed the DeepGauge modelling framework and, through rigorous theoretical treatment and in contrast to existing models, we proved that our framework gives estimates that satisfy all of the required theoretical properties of unit-level sets. In practice, this results in consistent inference of extremal dependence and provides a means to estimate joint tail properties; we demonstrate this by evaluating the performance of the DeepGauge framework on simulated and observed data in Sections 4 and 5, respectively. Unlike the majority of existing models for multivariate extremes, the DeepGauge framework is not limited to asymptotically dependent data nor low dimensions ($d\leq 4$). Thus, our approach represents a significant step towards flexible, non-parametric, and scalable models for multivariate extremes.
@@
One noticeable omission from the DeepGauge framework is a model for the angular density $f_{\bm{W}}(\bm{w}),\bm{w}\in\mathcal{S}^{d-1}$; see Section 3.4 for further discussion. Several existing approaches propose non-parametric techniques for this estimation, albeit in low dimensional ($d\leq 3$) settings (e.g., Papastathopoulos et al.,, 2024; Murphy-Barltrop et al., 2024a, ; Simpson and Tawn, 2024b, ). Estimates of the angular density allow one to simulate from the model described in equation (7), from which one can apply Monte Carlo techniques to obtain probability estimates for joint tail regions of any form, i.e., more general than those illustrated in Figure 4; see Wadsworth and Campbell, (2024) for instance. The estimation of, and simulation from, the angular density function $f_{\bm{W}}(\bm{w})$ therefore represents an important area for future work in the context of multivariate extremes.
@@
Declarations
Ethical Approval
Not Applicable
Availability of supporting data
The data sets analysed in Section 5 are available from the corresponding author upon reasonable request.
Competing interests
The authors have no relevant financial or non-financial interests to disclose.
Funding details
This work was supported by grants from the United States Geological Survey’s National Climate Adaptation Science Center (G24AC00197), and the National Science Foundation (DMS2152887).
Acknowledgments
The authors would like to thank Phil Jonathan for providing the data, Ryan Campbell for access to code, and Lambert de Monte for helpful discussions.
@@ A.1 Neural network representation
We here describe the construction of $r_{\tau}(\cdot)$ and $g(\cdot)$ used multi-layer perceptrons. Let $r_{\tau}(\bm{w}):=\exp\{m_{\bm{\psi}}(\bm{w})\},$ where $m_{\bm{\psi}}(\bm{w})$ is a neural network, parameterised by $\bm{\psi}$, with $N\in\mathbb{N}$ hidden layers. Then $m_{\bm{\psi}}$ is the composition ${m_{\bm{\psi}}(\cdot)=m_{\bm{\psi}}^{(N+1)}\circ\dots\circ\mathbf{m}^{(2)}_{\bm{\psi}}\circ\mathbf{m}^{(1)}_{\bm{\psi}}(\cdot)}$, where, for $j=1,\dots,N$, the output from the $j$-th hidden layer, $\mathbf{m}^{(j)}_{\bm{\psi}}:\mathbb{R}^{n_{j-1}}\mapsto\mathbb{R}^{n_{j}}$, can be written as
@@ A.2 Training and pre-training
The minimisation in (8) is performed using the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba,, 2014), which is a variant of mini-batch stochastic gradient descent. The Adam algorithm is applied with its hyper-parameters at their default settings, and the models are trained using the R interface to keras (Allaire and Chollet,, 2021). Prior to training, we uniformly-at-random partition the data into $80\%$ training and $20\%$ validation, with this partition consistent across the training stage for both $r_{\tau}(\bm{w})$ and $\tilde{g}(\bm{w})$. To mitigate overfitting of the neural network model, we train with checkpoints and early-stopping (Prechelt,, 2002). Neural network models are optimised for a finite number of pre-specified iterations, say $M$, by minimising the loss function evaluated on the training data. The validation data are not used to optimise the neural network parameters. Instead, at every iteration of the training algorithm, we evaluate the loss function on the validation data. Then, the “best fitting” model is determined to be that which which minimises the loss function, evaluated on the validation data, across all epochs. Early-stopping ensures that the training scheme does not necessarily run for all $M$ iterations; training stops early if the validation loss has not decreased in the last $\Delta\in\mathbb{N}$ iterations. We also add $L_{1}$ and $L_{2}$ penalties to the estimable neural network parameters, $\psi$, to provide further model regularisation; throughout we adopt these penalties with equal shrinkage weight of $1\times 10^{-4}$. For further details on fitting and regularising deep neural network models, see Goodfellow et al., (2016).
@@
Typically, without prior knowledge of the function that one seeks to approximate, the estimable parameter set, $\bm{\psi},$ of a neural network model is randomly initialised prior to training; thus when training our neural network model for the exceedance threshold $r_{\tau}(\cdot)$ we always choose to take this approach, as we have no prior knowledge of its structure. However, for training of the neural network that comprises the gauge function $g(\cdot)$, we can exploit some of the theoretical structure of gauge functions to perform pre-training, which is a popular technique used in applications of neural networks to reduce computation times. Pre-training uses parameter estimates of a neural network trained for one task as initial parameter estimates for a neural network model designed for a different task (with equivalent architecture; see Goodfellow et al., (2016), Ch. 8). Following Corollary 2.2, we can use an estimate of the radial exceedance threshold, say $\hat{r}_{\tau}(\bm{w}),$ to obtain an initial estimate for the gauge function, say $\hat{g}_{\tau}(\bm{w})$; this is achieved by replacing the radial function $h(\cdot)$ in Corollary 2.2 with $\hat{r}_{\tau}(\cdot)$ and deriving the corresponding rescaled gauge function. In the two dimensional setting, Wadsworth and Campbell, (2024) use a similar idea to get an initial estimate for the unit-level set $\partial\mathcal{G}$. The authors also show that, for $\tau$ close to 1, the rescaled quantile set is a valid approximation of $\partial\mathcal{G}$. We use this idea to pre-train the neural network of the gauge function model $\tilde{g}(\cdot)$; we perform an initial optimisation of its weights subject to minimisation of the squared-error loss between the initial gauge function $\hat{g}_{\tau}(\bm{w})$ and the model output, $\tilde{g}(\bm{w})$. For $\alpha$, we always set its initial estimate to $d$, as this is the theoretical value attained by gauge functions for many popular copula models (Wadsworth and Campbell,, 2024). In unreported experiments, we found that this pre-training helps to mitigate numerical instability during training and improves accuracy of gauge function estimation.
@@ Appendix B Additional proofs
is standard Laplace distributed, where the $\log(2)$ here denotes the median of $X_{E}$. Given $\bm{w}\in\mathcal{S}^{d-1}_{+}$ we have, by definition, $g_{E}(\bm{w})=\lim_{t\to\infty}(-\log f_{\bm{X}_{E}}(t\bm{w})/t)$, where $f_{\bm{X}_{E}}(\cdot)$ denotes the continuous joint density function for $\bm{X}_{E}$ (Nolde and Wadsworth,, 2022). For any large $t$ satisfying $t>\max_{i=1,\dots,d}\{\log(2)/w_{i}\}>0$, we have that
@@
Given any large $u$ satisfying $u>\max_{i=1,\dots,d}\{\log(2)/w_{i}\}>0$, and applying similar reasoning as in the proof of Proposition 2.1, we have that
@@
Given $\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$, set $\bm{c}:=\text{sgn}(\bm{w})=(\text{sgn}(w_{1}),\ldots,\text{sgn}(w_{d}))^{T}$. Then, we have that
@@
By Proposition 2.5, $g_{\bm{c}\bm{X}}(\bm{w}^{*})=g(\bm{w}^{*}/\bm{c})$, with $\bm{w}^{*}/\bm{c}:=(w^{*}_{1}/c_{1},\ldots,w^{*}_{d}/c_{d})^{T}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$ and $\text{sgn}(\bm{w})=\text{sgn}(\bm{w}^{*}/\bm{c})$. Thus, we have
@@
$\frac{\tilde{\mathfrak{r}}_{\bm{w}}}{||\bm{w}||_{\infty}}\geq\min_{i=1,\dots,d}\left\{\frac{|w_{i}|}{g(\bm{w})|w_{i}|}\right\}=\frac{1}{g(\bm{w})}.$
Recalling from Proposition 2.8 that $||\bm{w}||_{\infty}\Lambda(\bm{w})^{-1}=\tilde{\mathfrak{r}}_{\bm{w}}$, the first inequality follows. The second inequality follows directly from Proposition 2.8. ∎